Matt boasted about the size of the marlin he caught. "The head alone was 14 inches! The tail was the size of the head and half the body. The body was as long as the tail and head combined." What was the size of Matt's marlin?
Answer by edjones(8007)

t=14+ b/2
b=14+t
b=14+14+ b/2 Substitute 14+ b/2
b=28+ b/2
2b=56+b multiply each side by 2.
b=56
t=14+28=42
14+56+28=98" the size of Matt's marlin.
.
Ed